OpenStudy (anonymous):

six times the sum of a number and 3 is 12 less than 12 times the number . Write and solve an equation to find the number

OpenStudy (cwrw238):

6(x + 3) = 12x - 12 6x + 18 = 12x - 12 30 = 6x x = 30/6 = 5